The Ka of propanoic acid (C2H5COOH) is 1.34 × 10-5. Calculate the pH of the solution and the concentrations of C2H5COOH and C2H5COO– in a 0.421 M propanoic acid solution at equilibrium.
Answer #1

construct ICE table

C2H5COOH (aq) + H2O (l) <------------> C2H5COO- (aq) + H3O+ (aq)

I 0.421 0 0

C -x +x +x

E 0.421 - x +x +x

Ka = [C2H5COO-] [H3O+] / [C2H5COOH]

1.34 x 10^-5 = [x] [x] / [0.421-x]

x2 + x 1.34 * 10^-5 - 5.6414 * 10^-6 = 0

solve the quadratic equation

x = 0.002368 M

so concentrations of

[C2H5COO-] = x = 0.002368 M

[H3O+] = x = 0.002368 M

[C2H5COOH] = 0.421 - 0.002368 = 0.4186 M

pH = -log[H3O+]

= -log[0.002368]

= 2.625